OpenStudy (anonymous):

Rationalize the denominator of sqrt-9/(4-7i)-(6-6i)

OpenStudy (anonymous):

\[\frac{\sqrt{-9}}{(4-7i)-(6-6i)}=\frac{3i}{-2-i}*\frac{-2+i}{-2+i}=\frac{-6i-3}{4-i^2}=\frac{-6i-3}{5}\] I'm assuming that is how you meant to write your expression based on some previous questions...